OpenStudy (anonymous):

help please :) Find | |U + V| | given the vectors U = <0, 4> and V = <3, -5>. a) b) 3 c) 10 d) 2

OpenStudy (anonymous):

:)

OpenStudy (mathmate):

U+V=<0,4>+<3,-5> = <0+3,4-5>=<3,-1> None of the answers is suitable for ||U+V|| (magnitude) unless (c) reads sqrt(10).

OpenStudy (anonymous):

crap sorry. wrong one

OpenStudy (anonymous):

Find the angle between the vectors U = <-2, -1> and V = <-3, 1>. a) 58° b) 135° c) 45° d) 35°

OpenStudy (mathmate):

Do you know dot products?

OpenStudy (anonymous):

yes

OpenStudy (mathmate):

Do you know how to get the cosine between two vectors, given the two vectors?

OpenStudy (anonymous):

i do not think so

OpenStudy (mathmate):

How about finding unit vector from a vector?

OpenStudy (anonymous):

could you just give me the answer? haha. i konw the steps but i dont really know what they're called

OpenStudy (mathmate):

We're not supposed to directly give the answers. There's Wolfram which will give you answers blindly. You come here because you'd like to know how to work out the problem, unless I am mistaken.

OpenStudy (anonymous):

yeah, you're right. okay, just tell me how to set it up

OpenStudy (mathmate):

Basically, you would find the unit vector corresponding to each vector, say u for U, and v for V. Then the dot product (u.v) is the cosine of the angle between them. Try and see if you can work it out.

OpenStudy (anonymous):

im struggling with this.

OpenStudy (mathmate):

To find the unit vector, you would divide each element by the magnitude. U=<-2,-1>, and ||U||=sqrt((-2)^2+(-1)^2)=sqrt(5) So unit vector u=<-2/sqrt(5), -1/sqrt(5)>, or for simplicity, u=(1/sqrt(5))<-2,-1> Do the same for V. Find the dot product u.v and simplify to get cosine(theta).

OpenStudy (mathmate):

You can post a reply for each step if it is too long to do in one shot.

OpenStudy (anonymous):

okay i got 45

OpenStudy (anonymous):

hopefully thats right

OpenStudy (zzr0ck3r):

a dot b = |a||b|cos(theta) so cos^(-1)(a dot b)/(|a||b|) = theta

OpenStudy (mathmate):

That's what I got too!

OpenStudy (anonymous):

okay cool! thanks guys

OpenStudy (mathmate):

You're welcome! :)